# Three quantities arising from Bézout’s identity and resultants of integer polynomials

### v1

@@ Abstract.
In this paper, we study three quantities arising naturally from Bézout’s identity, the resultant and the reduced resultant of two non-zero coprime integer polynomials. We establish several new divisibility relations among them. We also pose two conjectures by making computations.
@@ 1. Introduction
In this paper, integer polynomials are meant to be polynomials in $\mathbb{Z}[x]$, that is, polynomials in the variable $x$ and with integer coefficients. They are definitely among the fundamental objects in algebra and number theory. The research around their arithmetic properties has a very long history and is still very active. The arithmetic properties include reducibility, degeneracy, decomposability, moduli of roots, signature of roots, multiplicative dependence of roots, etc. See [1, 3, 4, 5, 6, 9, 10, 11, 12, 13, 14, 15, 20] for some recent developments.
In this paper, we study the relations among three quantities arising naturally from Bézout’s identity, the resultant and the reduced resultant of two non-zero coprime integer polynomials. Here, we say two integer polynomials are coprime if they are coprime in the polynomial ring $\mathbb{Q}[x]$ over the rational numbers $\mathbb{Q}$.
From now on, $f(x)$ and $g(x)$ are two non-zero coprime integer polynomials. Below we define the three quantities studied in this paper with respect to $f(x)$ and $g(x)$.
By Bézout’s identity, we know that there exist uniquely two polynomials $u(x)$ and $v(x)$ in $\mathbb{Q}[x]$ such that:
(1.1) $u(x)f(x)+v(x)g(x)=1,\quad\deg(u)<\deg(g),\quad\deg(v)<\deg(f).$
Then, we define the quantity $B(f,g)$ to be the least common multiple of the denominators of all the coefficients of $u(x)$ and $v(x)$ in (1.1) (the coefficients are reduced to their lowest terms). Letting $B=B(f,g)$, we have
(1.2) $Bu(x)f(x)+Bv(x)g(x)=B,\quad Bu(x),Bv(x)\in\mathbb{Z}[x].$
Since $f(x)$ and $g(x)$ are non-zero and coprime, their resultant $\mathrm{Res}(f,g)$ is a non-zero integer. Then, we define $R(f,g)=|\mathrm{Res}(f,g)|$. It is well known that the resultant $\mathrm{Res}(f,g)$ can be expressed as a linear combination of $f(x)$ and $g(x)$ over the polynomial ring $\mathbb{Z}[x]$; see [7, page 157, Proposition 9]. So, there are two integer polynomials $s(x),t(x)$ such that
(1.3) $s(x)f(x)+t(x)g(x)=R(f,g).$
In addition, we define the quantity $r(f,g)$ to be the smallest positive integer in the following set
$f(x)\mathbb{Z}[x]+g(x)\mathbb{Z}[x]=\left\{u(x)f(x)+v(x)g(x):\,u(x),v(x)\in\mathbb{Z}[x]\right\}.$
This quantity is called reduced resultant in [22]. In addition, it is also called congruence number in [23, Definition 2.6], and moreover, the authors in [23] also described its applications.
@@
This is known to the experts.
@@
From the discussions in MathOverflow (see, for instance, [19, 24]), one can see that many people are interested with the relations among the three quantities $B(f,g)$, $r(f,g)$ and $R(f,g)$. However, so far we haven’t found a paper studying their relations systematically.
In this paper, we want to investigate systematically the divisibility relations among the three quantities $B(f,g)$, $r(f,g)$ and $R(f,g)$. We believe that some of them are new; such as Theorems 3.1, 4.1, 5.1 and 5.6. We also pose two conjectures at the end.
2. Preliminaries
In this section, we present some notation and some results which are needed later on.
First, we recall that for an integer polynomial $h(x)$, the content of $h$, denoted by $\mathrm{cont}(h)$, is the greatest common divisor of the coefficients of $h$. Besides, we denote by $L(h)$ the leading coefficient of $h$.
For two integer polynomials $h(x)$ and $q(x)$, recall that their resultant $\mathrm{Res}(h,q)$ is defined to be the determinant of their Sylvester matrix. We first list three classical properties of resultants as follows (see [7] or [17] for their proofs):
(2.1) $\mathrm{Res}(q,h)=(-1)^{\deg(h)\deg(q)}\mathrm{Res}(h,q),$
(2.2) $\mathrm{Res}(h,q_{1}q_{2})=\mathrm{Res}(h,q_{1})\mathrm{Res}(h,q_{2}),$
@@ 2. Preliminaries
where $h,q,q_{1},q_{2}$ are integer polynomials, and $c$ is a non-zero constant. The result in [8, Lemma 4.1] asserts that if we can write $h=sq+t$ with integer polynomials $s,t$, then
(2.4) $\mathrm{Res}(q,h)=L(q)^{\deg(h)-\deg(t)}\mathrm{Res}(q,t).$
For two non-zero coprime integer polynmials $f(x),g(x)$, we have known that $R(f,g)$ can be expressed as a linear combination of $f$ and $g$ in $\mathbb{Z}[x]$ (see (1.3)). The following result gives a more restricted expression of $R(f,g)$.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, let $d$ be the greatest common divisor of their leading coefficients. Then, there are two unique polynomials $u(x)$ and $v(x)$ such that $\deg(u)<\deg(g)$, $\deg(v)<\deg(f)$, $d\mid\mathrm{cont}(u)$, $d\mid\mathrm{cont}(v)$, and
$u(x)f(x)+v(x)g(x)=R(f,g).$
@@
$f(x)=a_{n}x^{n}+a_{n-1}x^{n-1}+\cdots+a_{1}x+a_{0},$
$g(x)=b_{m}x^{m}+b_{m-1}x^{m-1}+\cdots+b_{1}x+b_{0},$
with $a_{n}b_{m}\neq 0$. Regarding the Sylvester matrix of $f(x)$ and $g(x)$, for each $i$ ($11\leq i\leq n+m-$ 1), multiply the $i$-th column by $x^{n+m-i}$ and add it to the $(n+m)$-th column (here we view $x$ as an arbitrary non-zero integer). Then, we obtain
$\mathrm{Res}(f,g)=\det\left(\begin{array}[]{ccccccc}a_{n}&a_{n-1}&a_{n-2}&\cdots&0&0&x^{m-1}f\\ 0&a_{n}&a_{n-1}&\cdots&0&0&x^{m-2}f\\ 0&0&a_{n}&\cdots&0&0&x^{m-3}f\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&a_{1}&a_{0}&xf\\ 0&0&0&\cdots&a_{2}&a_{1}&f\\ b_{m}&b_{m-1}&b_{m-2}&\cdots&0&0&x^{n-1}g\\ 0&b_{m}&b_{m-1}&\cdots&0&0&x^{n-2}g\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&b_{1}&b_{0}&xg\\ 0&0&0&\cdots&b_{2}&b_{1}&g\end{array}\right).$
By the definition of $d$, we can write $a_{n}=ad,b_{m}=bd$ for two coprime positive integers $a,b$. Then, we obtain
$\mathrm{Res}(f,g)$ $=d\cdot\det\left(\begin{array}[]{ccccc}a&a_{n-1}&\cdots&0&x^{m-1}f\\ 0&a_{n}&\cdots&0&x^{m-2}f\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&a_{1}&f\\ b&b_{m-1}&\cdots&0&x^{n-1}g\\ 0&b_{m}&\cdots&0&x^{n-2}g\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&b_{1}&g\end{array}\right).$
@@
$\mathrm{Res}(f,g)$ $=d\Big{(}x^{m-1}fA_{1,m+n}+\cdots+fA_{m,m+n}+$
$\qquad\qquad x^{n-1}gA_{m+1,m+n}+\cdots+gA_{m+n,m+n}\Big{)}$
$=d(x^{m-1}A_{1,m+n}+\cdots+A_{m,m+n})f+$
$\qquad\qquad d(x^{n-1}A_{m+1,m+n}+\cdots+A_{m+n,m+n})g$
$=u_{1}(x)f(x)+v_{1}(x)g(x),$
@@
$u_{1}(x)=d(x^{m-1}A_{1,m+n}+...+A_{m,m+n})\in\mathbb{Z}[x],$
$v_{1}(x)=d(x^{n-1}A_{m+1,m+n}+\cdots+A_{m+n,m+n})\in\mathbb{Z}[x].$
@@
(2.5) $R(f,g)=|\mathrm{Res}(f,g)|=u(x)f(x)+v(x)g(x),$
where $u(x)=u_{1}(x)$ and $v(x)=v_{1}(x)$ if $\mathrm{Res}(f,g)>0$, and otherwise $u(x)=-u_{1}(x)$ and $v(x)=-v_{1}(x)$.
Note that in the above deductions, we view $x$ as an arbitrary integer. That is, (2.5) holds for any integer $x$. So, (2.5) also holds if we view $x$ as a variable.
@@
$d\mid\mathrm{cont}(u),\qquad d\mid\mathrm{cont}(v).$
Meanwhile, it can be seen that $\deg(u)\leq m-1<\deg(g)$, $\deg(v)\leq n-1<\deg(f)$. Finally, since $f$ and $g$ are coprime, one can prove by contradiction that $u(x)$ and $v(x)$ are unique. ∎
@@ 3. Relations between $B(f,g)$ and $r(f,g)$
In this section, we investigate the relation between $B(f,g)$ and $r(f,g)$ for two non-zero coprime integer polynomials $f(x),g(x)$. We have already known that $r(f,g)\mid B(f,g)$ by (1.4). The following result is about the converse of this divisibility relation.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, there exists a non-negative integer $k$ such that
$B(f,g)\mid d^{k}r(f,g),$
where $d$ is the greatest common divisor of the leading coefficients of $f(x)$ and $g(x)$.
@@
For simplicity, let $B=B(f,g)$ and $r=r(f,g)$. By definition, there exist two integer polynomials $u(x)$ and $v(x)$ such that
(3.1) $u(x)f(x)+v(x)g(x)=r.$
@@
(3.2) $\deg(u)+\deg(f)=\deg(v)+\deg(g).$
So, either $\deg(u)<\deg(g)$ and $\deg(v)<\deg(f)$, or $\deg(u)\geq\deg(g)$ and $\deg(v)\geq\deg(f)$.
If $\deg(u)<\deg(g)$ and $\deg(v)<\deg(f)$, then we have
$\frac{u(x)}{r}f(x)+\frac{v(x)}{r}g(x)=1.$
@@
Now, we assume that $\deg(u)\geq\deg(g)$ and $\deg(v)\geq\deg(f)$. Recall that $L(f)$ and $L(g)$ are the leading coefficients of $f$ and $g$ respectively. Let $d=\gcd(L(f),L(g))$. From (3.1) and noticing $r\in\mathbb{Z}$, we have
$L(u)\cdot L(f)=-L(v)\cdot L(g),$
where $L(u)$ and $L(v)$ are the leading coefficients of $u$ and $v$ respectively. Moreover, we get
(3.3) $L(u)\cdot\frac{L(f)}{d}=-L(v)\cdot\frac{L(g)}{d}.$
Since $\gcd(\frac{L(f)}{d},\frac{L(g)}{d})=1$, we have $\frac{L(f)}{d}\mid L(v)$ and $\frac{L(g)}{d}\mid L(u)$. Then, we define two integer polynomials:
$h_{1}(x)=\frac{dL(u)}{L(g)}x^{\deg(u)-\deg(g)},\qquad q_{1}(x)=\frac{dL(v)}{L(f)}x^{\deg(v)-\deg(f)}.$
@@
$d\cdot u(x)=g(x)h_{1}(x)+u_{1}(x),$
$d\cdot v(x)=f(x)q_{1}(x)+v_{1}(x),$
where $u_{1}(x)$ and $v_{1}(x)$ are some integer polynomials such that $\deg(u_{1})<\deg(u)$ and $\deg(v_{1})<\deg(v)$. Combining the above two formulas with (3.1), we get
$f(x)g(x)(h_{1}(x)+q_{1}(x))+u_{1}(x)f(x)+v_{1}(x)g(x)=dr,$
@@
$u_{1}(x)f(x)+v_{1}(x)g(x)=dr.$
Repeating the above steps, after a finite number of steps (say $k$ steps), there must exist integer polynomials $u_{k}(x)$ and $v_{k}(x)$ such that $\deg(u_{k})<\deg(g)$ and $\deg(v_{k})<\deg(f)$, and
$u_{k}(x)f(x)+v_{k}(x)g(x)=d^{k}r.$
@@
We remark that the non-negative integer $k$ in Theorem 3.1 is at most $\max(\deg(u)-\deg(g)+1,0)$ with polynomial $u$ in (3.1). Moreover, one can find an upper bound for $\deg(u)$ in [2, Theorem A].
When the leading coefficients of $f(x)$ and $g(x)$ are coprime (that is, $d=1$ in Theorem 3.1), we have $B(f,g)\mid r(f,g)$. Then, noticing $r(f,g)\mid B(f,g)$, we directly get the following corollary.
@@
For any two non-zero coprime integer polynomials $f(x)$ and $g(x)$, if their leading coefficients are coprime, then $B(f,g)=r(f,g)$.
If the leading coefficients of $f(x)$ and $g(x)$ are not coprime, then either $B(f,g)=r(f,g)$ or $B(f,g)\neq r(f,g)$. Both cases can happen; see Examples 3.3 and 3.4.
@@
We also remark that if the leading coefficients of $f(x)$ and $g(x)$ are not coprime and let $d$ be their greatest common divisor, then $B(f,g)$ and $r(f,g)$ may not have the same prime factors, and also $B(f,g)$ and $dr(f,g)$ may not have the same prime factors; see Example 3.4. However, by Theorem 3.1 we know that any prime factor of $B(f,g)$ divides $dr(f,g)$.

### v2

# Three integers arising from Bézout’s identity and resultants of integer polynomials

@@ Abstract.
In this paper, we study three integers arising naturally from Bézout’s identity, the resultant and the reduced resultant of two coprime integer polynomials. We establish several new divisibility relations among them. We also pose two conjectures by making computations.
@@ 1. Introduction
In this paper, integer polynomials are meant to be polynomials in $\mathbb{Z}[x]$, that is, polynomials in the variable $x$ and with integer coefficients. They are definitely among the fundamental objects in algebra and number theory. The research around their arithmetic properties has a very long history and is still very active. The arithmetic properties include reducibility, degeneracy, decomposability, moduli of roots, signature of roots, multiplicative dependence of roots, etc. See [1, 4, 5, 6, 7, 8, 11, 12, 13, 14, 15, 16, 17, 18, 19, 22, 25, 32] for some recent developments.
In this paper, we study the relations among three integers arising naturally from Bézout’s identity, the resultant and the reduced resultant of two coprime integer polynomials with positive degree. Here, we say two integer polynomials are coprime if they don’t have a common root.
From now on, $f(x)$ and $g(x)$ are two coprime integer polynomials of positive degree. Below we define the three integers studied in this paper with respect to $f(x)$ and $g(x)$.
By Bézout’s identity, we know that there exist uniquely two polynomials $p(x)$ and $q(x)$ in $\mathbb{Q}[x]$ such that:
(1.1) $p(x)f(x)+q(x)g(x)=1,\quad\deg(p)<\deg(g),\quad\deg(q)<\deg(f).$
Then, we define the integer $B(f,g)$ to be the least common multiple of the denominators of all the coefficients of $p(x)$ and $q(x)$ in (1.1) (the coefficients are reduced to their lowest terms). Letting $B=B(f,g)$, we have
(1.2) $Bp(x)f(x)+Bq(x)g(x)=B,\quad Bp(x),Bq(x)\in\mathbb{Z}[x].$
Since $f(x)$ and $g(x)$ are non-constant and coprime, their resultant $\mathrm{Res}(f,g)$ is a non-zero integer. Then, we define the integer $R(f,g)=|\mathrm{Res}(f,g)|$. It is well known that the resultant $\mathrm{Res}(f,g)$ can be expressed as a linear combination of $f(x)$ and $g(x)$ over the polynomial ring $\mathbb{Z}[x]$; see [9, page 157, Proposition 9].
In addition, we define the integer $r(f,g)$ to be the smallest positive integer in the following set
$f(x)\mathbb{Z}[x]+g(x)\mathbb{Z}[x]=\left\{p(x)f(x)+q(x)g(x):\,p(x),q(x)\in\mathbb{Z}[x]\right\}.$
This integer is called reduced resultant in [28]. In addition, it is also called congruence number in [30, Definition 2.6], and moreover, the authors in [30] also described its applications.
@@
This was known previously.
@@
From the discussions in MathOverflow (see, for instance, [24, 29, 31]), one can see that many people are interested with the relations among the three integers $B(f,g)$, $r(f,g)$ and $R(f,g)$. However, so far we haven’t found a paper studying their relations systematically.
In this paper, we want to investigate systematically the divisibility relations among the three integers $B(f,g)$, $r(f,g)$ and $R(f,g)$. We believe that some of them are new; such as Theorems 3.1, 4.1, 5.1 and 5.6. We also pose two conjectures at the end.
To avoid systematic repetitions, throughout the paper we fix the following assumptions and notations:
  • $f(x)$ and $g(x)$ are two coprime integer polynomials of positive degree;
  • for every $h\in\mathbb{Z}[x]$, $L(h)$ is the leading coefficient of $h$, and $\mathrm{cont}(h)$ is the greatest common divisor of its coefficients;
  • $d=\gcd(L(f),L(g))$.
2. About the resultant
In this section, we present some results about the resultant $\mathrm{Res}(f,g)$ which are needed later on.
For two integer polynomials $h_{1}(x)$ and $h_{2}(x)$, recall that their resultant $\mathrm{Res}(h_{1},h_{2})$ is defined to be the determinant of their Sylvester matrix. We first list three classical properties of resultants as follows (see [9] or [21] for their proofs):
(2.1) $\mathrm{Res}(h_{1},h_{2})=(-1)^{\deg(h_{1})\deg(h_{2})}\mathrm{Res}(h_{2},h_{1}),$
(2.2) $\mathrm{Res}(h,h_{1}h_{2})=\mathrm{Res}(h,h_{1})\mathrm{Res}(h,h_{2}),$
@@ 2. About the resultant
where $h,h_{1},h_{2}$ are integer polynomials of positive degree, and $c$ is a non-zero constant. The result in [10, Lemma 4.1] asserts that if we can write $h_{2}=sh_{1}+t$ with integer polynomials $s,t$, then
(2.4) $\mathrm{Res}(h_{1},h_{2})=L(h_{1})^{\deg(h_{2})-\deg(t)}\mathrm{Res}(h_{1},t).$
The following result except the divisibility part $d\mid\mathrm{cont}(p)$, $d\mid\mathrm{cont}(q)$” is well-known (see [3, Proposition 4.18] or [26, Lemma 7.2.1]). We follow the approach in [26, Lemma 7.2.1] to recover the well-known part and also prove the divisibility part.
@@
There are two unique integer polynomials $p(x)$ and $q(x)$ such that $\deg(p)<\deg(g)$, $\deg(q)<\deg(f)$, $d\mid\mathrm{cont}(p)$, $d\mid\mathrm{cont}(q)$, and
$p(x)f(x)+q(x)g(x)=R(f,g).$
@@
$f(x)=a_{m}x^{m}+a_{m-1}x^{m-1}+\cdots+a_{1}x+a_{0},$
$g(x)=b_{n}x^{n}+b_{n-1}x^{n-1}+\cdots+b_{1}x+b_{0},$
with $a_{m}b_{n}\neq 0$. Regarding the Sylvester matrix of $f(x)$ and $g(x)$, for each $i$ ($1\leq i\leq m+n-1$), we multiply the $i$-th column by $x^{m+n-i}$ and add it to the $(m+n)$-th column (here we view $x$ as an arbitrary non-zero integer). Then, we obtain
$\mathrm{Res}(f,g)=\det\left(\begin{array}[]{ccccccc}a_{m}&a_{m-1}&a_{m-2}&\cdots&0&0&x^{n-1}f\\ 0&a_{m}&a_{m-1}&\cdots&0&0&x^{n-2}f\\ 0&0&a_{m}&\cdots&0&0&x^{n-3}f\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&a_{1}&a_{0}&xf\\ 0&0&0&\cdots&a_{2}&a_{1}&f\\ b_{n}&b_{n-1}&b_{n-2}&\cdots&0&0&x^{m-1}g\\ 0&b_{n}&b_{n-1}&\cdots&0&0&x^{m-2}g\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\cdots&b_{1}&b_{0}&xg\\ 0&0&0&\cdots&b_{2}&b_{1}&g\end{array}\right).$
In view of $d=\gcd(L(f),L(g))$, we can write $a_{m}=ad,b_{n}=bd$ for two coprime integers $a,b$. Then, we obtain
$\mathrm{Res}(f,g)$ $=d\cdot\det\left(\begin{array}[]{ccccc}a&a_{m-1}&\cdots&0&x^{n-1}f\\ 0&a_{m}&\cdots&0&x^{n-2}f\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&a_{1}&f\\ b&b_{n-1}&\cdots&0&x^{m-1}g\\ 0&b_{n}&\cdots&0&x^{m-2}g\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&\cdots&b_{1}&g\end{array}\right).$
@@
$\mathrm{Res}(f,g)$ $=d\Big(x^{n-1}fA_{1,m+n}+\cdots+fA_{n,m+n}+$
$\qquad\qquad x^{m-1}gA_{n+1,m+n}+\cdots+gA_{m+n,m+n}\Big)$
$=d(x^{n-1}A_{1,m+n}+\cdots+A_{n,m+n})f+$
$\qquad\qquad d(x^{m-1}A_{n+1,m+n}+\cdots+A_{m+n,m+n})g$
$=p_{1}(x)f(x)+q_{1}(x)g(x),$
@@
$p_{1}(x)=d(A_{1,m+n}x^{n-1}+\cdots+A_{n,m+n})\in\mathbb{Z}[x],$
$q_{1}(x)=d(A_{n+1,m+n}x^{m-1}+\cdots+A_{m+n,m+n})\in\mathbb{Z}[x].$
@@
(2.5) $R(f,g)=|\mathrm{Res}(f,g)|=p(x)f(x)+q(x)g(x),$
where $p(x)=p_{1}(x)$ and $q(x)=q_{1}(x)$ if $\mathrm{Res}(f,g)>0$, and otherwise $p(x)=-p_{1}(x)$ and $q(x)=-q_{1}(x)$.
Note that in the above deductions, we view $x$ as an arbitrary non-zero integer. That is, (2.5) holds for any non-zero integer $x$. So, (2.5) also holds if we view $x$ as a variable.
@@
$d\mid\mathrm{cont}(p),\qquad d\mid\mathrm{cont}(q).$
Meanwhile, it can be seen that $\deg(p)\leq m-1<\deg(g)$, $\deg(q)\leq n-1<\deg(f)$. Finally, since $f$ and $g$ are coprime, one can prove by contradiction that $p(x)$ and $q(x)$ are unique. ∎
@@ 3. Relations between $B(f,g)$ and $r(f,g)$
We have already known that $r(f,g)\mid B(f,g)$ by (1.3). The following result is about the converse of this divisibility relation. Recall that $d=\gcd(L(f),L(g))$.
@@
There exists a non-negative integer $k$ such that
$B(f,g)\mid d^{k}r(f,g).$
@@
For simplicity, let $B=B(f,g)$ and $r=r(f,g)$. By definition, there exist two integer polynomials $p(x)$ and $q(x)$ such that
(3.1) $p(x)f(x)+q(x)g(x)=r.$
@@
(3.2) $\deg(p)+\deg(f)=\deg(q)+\deg(g).$
So, either $\deg(p)<\deg(g)$ and $\deg(q)<\deg(f)$, or $\deg(p)\geq\deg(g)$ and $\deg(q)\geq\deg(f)$.
If $\deg(p)<\deg(g)$ and $\deg(q)<\deg(f)$, then we have
$\frac{p(x)}{r}f(x)+\frac{q(x)}{r}g(x)=1.$
@@
Now, we assume that $\deg(p)\geq\deg(g)$ and $\deg(q)\geq\deg(f)$. From (3.1) and noticing $r\in\mathbb{Z}$, we have L(pf)=L(qg)L(pf)=-L(qg), that is,
$L(p)\cdot L(f)=-L(q)\cdot L(g).$
Moreover, we get
(3.3) $L(p)\cdot\frac{L(f)}{d}=-L(q)\cdot\frac{L(g)}{d}.$
Since $\gcd(\frac{L(f)}{d},\frac{L(g)}{d})=1$, we have $\frac{L(f)}{d}\mid L(q)$ and $\frac{L(g)}{d}\mid L(p)$. Then, we define two integer polynomials:
$h(x)=\frac{dL(p)}{L(g)}x^{\deg(p)-\deg(g)},\qquad t(x)=\frac{dL(q)}{L(f)}x^{\deg(q)-\deg(f)}.$
@@
$d\cdot p(x)=g(x)h(x)+p_{1}(x),$
$d\cdot q(x)=f(x)t(x)+q_{1}(x),$
where $p_{1}(x)$ and $q_{1}(x)$ are some integer polynomials such that $\deg(p_{1})<\deg(p)$ and $\deg(q_{1})<\deg(q)$. Combining the above two formulas with (3.1), we get
$f(x)g(x)(h(x)+t(x))+p_{1}(x)f(x)+q_{1}(x)g(x)=dr,$
@@
$p_{1}(x)f(x)+q_{1}(x)g(x)=dr.$
Repeating the above steps, after a finite number of steps (say $k$ steps), there must exist integer polynomials $p_{k}(x)$ and $q_{k}(x)$ such that $\deg(p_{k})<\deg(g)$, $\deg(q_{k})<\deg(f)$, and
$p_{k}(x)f(x)+q_{k}(x)g(x)=d^{k}r.$
@@
We remark that the non-negative integer $k$ in Theorem 3.1 is at most $\max(\deg(p)-\deg(g)+1,0)$ with polynomial $p$ in (3.1). So, in Theorem 3.1 we can choose
$k=\max(\deg(p)-\deg(g)+1,0).$
Moreover, one can find an upper bound on $\deg(p)$ for one choice of $p(x)$ in [2, Theorem A].
When $d=1$ in Theorem 3.1, we have $B(f,g)\mid r(f,g)$. Then, noticing $r(f,g)\mid B(f,g)$, we directly get the following corollary. We believe that this result is known to the experts (see, for instance, Myerson’s answer in [29] for the case of monic polynomials), but we haven’t found a specific reference for it.
@@
If $d=1$, then $B(f,g)=r(f,g)$.
If $d>1$, then either $B(f,g)=r(f,g)$ or $B(f,g)\neq r(f,g)$. Both cases can happen; see Examples 3.3 and 3.4.
@@
We also remark that if $d>1$, then $B(f,g)$ and $r(f,g)$ may not have the same prime factors, and also $B(f,g)$ and $dr(f,g)$ may not have the same prime factors; see Example 3.4. However, by Theorem 3.1 we know that any prime factor of $B(f,g)$ divides $dr(f,g)$.